A free electron and a free proton are exactly 1.2cm apart.

A) Find the magnitude of the acceleration of the proton

B) Find the magnitude of the acceleration of the electron
Answer #1

F = kq1q2/r^2

F = 9x10^9*1.6x10^-19*1.6x10^-19/0.012^2 = 1.6x10^-24 N

a) acc_proton = F/m_p = 1.6x10^-24/1.67x10^-27 = 958.084 m/s^2

b) acc_ele = F/m_e = 1.6x10^24/9.1x10^-31 = 1.758x10^6 m/s